10, Ruiz Corporation is an unlevered Timber company with assets of $8.36 million. Ruiz reports earnings before interest and tax of $1.25 million. In order to save on taxes, the CFO has suggested that the firm issue debt and use the proceeds to retire shares of stock. If conducted, the capital structure change would result in $1.00 million of new debt with an annual interest expense of 12 percent.

a. How much will Ruiz save in taxes, per year, as a result of the decision to issue debt? Use a corporate tax rate of 40 percent.

Answer:$ ____million

Place your answer in millions with at least two decimal places, that is, if your answer is three quarters of a million dollars you should answer 0.75.

b.Suppose instead that the debt is permanent, meaning that the same amount of debt will stay on the books year after year forever. Under this scenario determine the how much in value the permanent debt tax shields provide?

Answer:$ ______million

Place your answer in millions with at least two decimal places, that is, if your answer is three quarters of a million dollars you should answer 0.75.
Answer #1
Particular Before debt (in million $) After debt (in million $)
EBIT 1.25 1.25
Less: Interest Expense 0 0.12
PBT 1.25 1.13
Less: Tax Expense 0.5 0.452

Interest Expense = Debt*Interest Rate

Tax Expense = Tase Rate*PBT

a.

Savings in taxes per year = 0.5-0.452 = $0.048 million

b.

Debt tax Shield = Tax Rate*Interest Expense = 0.40*0.12 = $0.048 million

Each year Debt tax shield is obtained, taking interest rate @12%

Value of total debt tax shield = 0.048/0.12 = $0.4 million